# Symmetry, Invariant Manifolds and Flow Reversals in Active Nematic Turbulence

### v1

@@ I Introduction
The remainder of the paper is organized as follows. In Section II, we introduce the system model, summarize our previous ECS-based analysis of active nematic channel flow, and review experimental and theoretical work on flow reversals in active and passive fluids. In Section III, we use the symmetries of the 2D AN equations to carry out an equivariant bifurcation analysis in minimal flow units, tracing primary and secondary local and global bifurcations, and constructing a symmetry-organized catalog of dynamically relevant ECSs. In Section IV, we use this catalog to describe transient flow reversals in the preturbulent regime, identifying a small number of ECS families and their heteroclinic connections as the main mechanisms by which trajectories switch between opposite nearly unidirectional states. In Section V, we extend this picture into the active turbulent regime in two MFUs, characterize the chaotic and quasiperiodic sets that arise at high activity, and demonstrate that persistent two-way reversals are organized by a reduced ECS skeleton whose members can be related back to the preturbulent branches via their symmetry signatures.
@@ II.1 Governing Equations
The domain is taken to be a two-dimensional channel of size $0,\tilde{L}]\times[0,h]$, with $x\in[0,\tilde{L}$] the streamwise (horizontal) direction and $y\in[0,h]$ the wall-normal (vertical) direction. We impose no-slip boundary conditions on the velocity field $\mathbf{u}$ and perpendicular (homeotropic) anchoring on the nematic director $\mathbf{n}$ at the confining walls. We employ the nondimenzionalized equations obtained in [27]:
@@ II.3 Flow Reversals in Active Nematics and Other Active Fluids
The Ph.D. thesis [39] reported experimental results on AN systems in long, narrow channels whose open ends were connected to an external chaotic AN reservoir. Persistent, nearly unidirectional flows were observed, and the direction of these uniaxial flow states (UNI) was selected at random via spontaneous symmetry breaking of the zero-flow (ZF) state, as expected. More importantly, spontaneous and repeated reversals were also observed between the left- and right-flowing UNI states, and the reversal frequency depended on the channel dimensions. The author used the channel-averaged streamwise velocity $\langle u\rangle$ as an order parameter to quantify these reversals. In these experiments, a typical reversal proceeded via an intermediate state: the UNI state with a parabolic velocity profile ($\langle u\rangle\approx\pm u_{0}$) first collapsed into a vortex-dominated configuration with negligible net flow ($\langle u\rangle\approx 0$), and then evolved into a UNI state whose mean flow was opposite in direction to the original one.
@@
In polar active fluids, recent experimental, computational and analytical work [40] has argued that flow reversals are a precursor to active turbulence. The authors found that a dense suspension of swimming Bacillus subtilis confined to a cylindrical wells exhbit a single vortex steady flow at small radii. This state, which is analogous to the UNI in channel flow, loses stability as the radius is increased beyond a threshold. In the resulting oscillatory state, the vortex periodically reverses its direction. Using simulations and weakly nonlinear analysis of a Toner–Tu–Swift–Hohenberg model, it was shown that the onset of reversals is controlled by a global infinite-period bifurcation mediated by a resonant nonlinear interaction of the three lowest modes of the flow.
@@ II.4 Possible Generic Dynamical Mechanisms Underlying Flow Reversals
Beyond active fluids, flow reversals have been extensively observed and studied in various settings [41, 42, 43, 44]. A prominent example is that of turbulent Rayleigh-Benard convection (RBC) [45], in which large-scale convection (LSC) reversals between the clockwise and anticlockwise vortical states are robustly observed [46]. In addition to numerical simulations, such reversals have been studied using both determninistic and stochastic reduced-order dynamical models [47, 42].
@@ III.2 Symmetries in the active nematics equations
Figure 1: The velocity field snapshots of some ECSs at representative activity numbers $\mathrm{R_{a}}=1.5$ and $\mathrm{R_{a}}=4.5$ are shown here with their symmetries in the full-length channel, $L=\tilde{L}$. The color map indicates vorticity.
@@ III.4.1 Traveling waves emerge from UNI
Figure 4: Projection onto the reduced phase space ($\langle Q_{11}\rangle$, $\langle v^{2}\rangle$), showing the stable and unstable TWs before the SNIPER bifurcation, with a heteroclinic connection passing close to UNI. Note the RPO that emerges after SNIPER has a shape similar to that of the vanished heteroclinic network.
@@ III.5.3 One-vortex standing waves via symmetry-preserving bifurcation
Figure 13: Two dancing-disclination-like POs (DPOs) that co-exist at $\mathrm{R_{a}}=1.5$, projected onto the reduced phase space ($\langle Q_{11}\rangle$$\langle v^{2}\rangle$). Both solutions lie in the $\langle u\rangle=0$ subspace (i.e., not net streamwise flow) due to the $\sigma_{x}\sigma_{y}$ symmetry.
@@ III.5.5 Secondary Bifurcations Associated with Higher Unstable Eigenvalues of LAN
So far, our analysis has been limited to bifurcations associated with the first unstable eigenspace of LAN. As the activity number is varied from $\mathrm{R}_{\mathrm{a}}\approx 0.85$ to $\mathrm{R}_{\mathrm{a}}\approx 4.5$, a total of eight (equal) pairs of real eigenvalues eventually cross the imaginary asix. The isotropy subgroups of these eight unstable 2D eigenspaces are listed in Table 4 in the Appendix. While the study of all bifurcations is outside the scope of this work, we briefly discuss the bifurcations from the second unstable eigenspace, since these solutions play an important role in the turbulent regime we analyze later in this paper.
@@
Figure 14: Visualization of POT2a at $\mathrm{R_{a}}=2.25$. (a) A snapshot of the nematic director and velocity fields, where its $\sigma_{x}\sigma_{y}\tau_{x}(L/2)$ symmetry is apparent. (b) Projection onto the reduced phase space: ($\langle Q_{11}\rangle$, $\langle v^{2}\rangle$). This PO lies in the $\langle u\rangle=0$ subspace (i.e., not net streamwise flow) due to the $\sigma_{x}\sigma_{y}$ symmetry.
@@ IV.2 Transient Flow Reversals at Higher Activity and the Path Toward Turbulence
The chaotic sets $\mathcal{X}$ and $\mathcal{Y}$ exhibit reversals between regions where the flow velocity has a mean value $\langle u\rangle$ that persistently oscillates about positive and negative values, respectively. In the following section, we analyze these attractors to elucidate the mechanisms responsible for flow reversals in the MUF at $\mathrm{R}_{\mathrm{a}}=4.5$.
@@ V Flow Reversal Mechanisms in Turbulence: Two-Way Persistent Reversals
We find that an arbitrary perturbation (i.e., with no symmetry) always ends up in $\mathcal{Y}$, indicating that in MFU A, $\mathcal{Y}$ is the only almost-global attractor while the other two sets $\mathcal{X}$ and $\mathcal{Z}$ are saddles. Motivated by the apparent $\tau_{x}(L_{A}/2)$ symmetry of $\mathcal{X}$ itself, we verify that in fact, all initial conditions with $\tau_{x}(L/2)$ symmetry end up in $\mathcal{X}$ in MFU A. Hence, we also analyze dynamics in the smaller MFU B, where $\mathcal{Y}$ is eliminated due to lack of two-fold translation symmetry, and we find that the halved-version of $\mathcal{X}$ is the only almost-global attractor that attracts arbitrary initial conditions almost surely. The quasiperiodic $\mathcal{Z}$, while still present in MFU B, again only attracts a measure zero set of trajectories that have the additional translation symmetry. Table 1 summarizes this discussion and answers Q1 posed above. Now we turn to answering Q2 and Q3.
@@ V.2.1 Dominant ECSs Within $\mathcal{X}$
(b) Reduced phase space projection showing the segment turbulent trajectory $X(t)$ in the interval $29000\tau\leq t\leq 30000\tau$ (black) and the six main ECS groups. The color scheme for the ECSs is the same as in (a). The orange circle indicates the beginning of the segment ($t=29000\tau$) and the magenta square indicates the end of the segment ($t=30000\tau$).
(c) Reduced phase space portrait showing the six main ECS groups shadowed by the turbulent trajectory, along with their symmetry-related copies under $\sigma_{y}$.
@@
Across the 2,501 points, 40 of the 75 ECS groups appear as the closest state at least once. To isolate the most dynamically significant structures, we focus on the ten most frequently encountered ECS groups. Together, these ten groups account for approximately 62% of all occurrences (see Table 6 in Appendix). As anticipated, each of these ten groups consists of ECSs that are invariant under $\tau_{x}(L_{A}/2)$ symmetry.
Figure 21: Symmetry-reduced distance plot showing the closest encounters between the chaotic trajectory $X(t)$ and the ten relevant ECS and their symmetry counterparts. The ECS list on the left follows the same color scheme as in Fig. 20(a). The plot shows a clear pattern of repeated transitions (i.e., reversals) between the ‘left flowing’ (blue) and ’right flowing’ (red) ECSs, via the intermediate (green) ECSs.
These ten ECS groups serve as a reduced representation that captures the essential skeleton of the attractor $\mathcal{X}$’s dynamics. Six of these groups belong to the green category and are natural candidates for mediating transitions between the red and blue regions, i.e., for enabling reversals. Figure 21 shows the distance plot between the trajectory and these ECS groups, highlighting the paths of reversals. In particular, the four ECS groups from the red/blue categories, along with those from the green category, are closely approximated by the trajectory. For example, in the interval $29000\tau,29200\tau]$, during the first reversal of the series (see also Fig. 20a), the distance plot shows that the trajectory first approaches the blue ECS group, with the mean velocity oscillating around $\langle u\rangle\approx-2$. It then nears some ECSs of the green group, where $\langle u\rangle\approx 0$, before finally transitioning to a state where the mean velocity oscillates around $\langle u\rangle\approx 2$ and closely follows the red ECS group.
@@ V.2.3 Reversal Mechanisms and Heteroclinic connections
First, the green ECSs play a role analogous to that of the ‘intermediary’ DPOs in the preturbulent regime. These green ECSs are likely originated from solutions invariant under symmetries conjugate to $\sigma_{x}\sigma_{y}$ and $\tau_{x}(L_{A}/2)$ in the preturbulent regime. These ECSs in the turbulent regime have slightly broken $\sigma_{x}\sigma_{y}$ symmetry, but retain the two-fold $\tau_{x}(L/2)$ translational symmetry. Supporting evidence includes their similarity to POT2a, and the observation that when a green solution is shadowed, its $\sigma_{x}\sigma_{y}$ counterpart is also shadowed (see Fig. 21) simultaneously, implying that these solutions with broken $\sigma_{x}\sigma_{y}$ symmetry lie close to each other. Second, the blue and red ECSs play a role analogous to that of HRPOs, and form the structural backbone for generating left- and right-flowing, nearly uniaxial states. Their phase space structures suggests that these ECSs are most likely the result of symmetry-breaking bifurcations of the HRPOs. The blue/red ECS, which retain the homoclinic-like structure in phase space, have broken the shift-reflect symmetries of their preturbulent counterparts, but still possess the two-fold $\tau_{x}(L_{A}/2)$ translation symmetry. Finally, the heteroclinic connections between these groups are likely the mechanisms through which turbulence exhibits flow reversals.
@@
Indeed, we observe that heteroclinic connections may exist between pairs of ECSs from different regions that are shadowed sequentially. For example, the red ECS $\sigma_{x}\sigma_{y}$T2/1 is shadowed during the interval $20067.20\tau,\,29073.6\tau]$, and immediately afterward, the green ECS $\sigma_{y}$T2/30 is shadowed during $29072.80\tau,\,29073.6\tau]$. This suggests a heteroclinic connection between these two ECSs. However, within the green region, the turbulent trajectory sequentially (and sometimes simultaneously) shadows several ECSs, including $\sigma_{y}$T2/30, $\sigma_{x}$T2/25, T2/13, $\sigma_{y}$T2/9, $\sigma_{x}$T2/30, $\sigma_{x}\sigma_{y}$T2/4, and $\sigma_{x}\sigma_{y}$T2/13. Consequently, identifying a heteroclinic connection to a target green ECS is challenging, as shadowing events at the red/green and blue/green boundaries involve multiple ECSs and are not sharply defined. Although we cannot conclusively verify a heteroclinic connection between $\sigma_{x}\sigma_{y}$T2/1 and $\sigma_{y}$T2/30, the turbulent trajectory shadows the unstable manifold of the former ECS and eventually approaches the latter. Throughout the subsequent exploration of the green region, the trajectory continues to follow the unstable manifold, as illustrated by the segment of $X(t)$ and a nearby trajectory in Fig. 26 within the interval $29064.0\tau,\,29094.4\tau]$.
@@
In this MFU B based calculation, we focus on the two most relevant blue/red ECSs and the three most relevant green ECSs (and their symmetry counterparts). In this smaller domain, the RPOs T2/1 and T2/15 each have two unstable directions, while the RPOs T2/9, T2/30, and T2/4 have three. We consider the dominant unstable direction to generate a sample of trajectories within each unstable manifold. Several of these trajectories pass nearby and shadow other ECSs, providing evidence of heteroclinic connections. Table 2 lists few of the pairs of ECSs connected by such trajectories, together with the corresponding minimum distances and shadowing time intervals. The Figs. 27a and 27b highlight two such examples.

### v2

@@ I Introduction
The remainder of this paper is organized as follows. In Section II, we introduce the system model, summarize our previous ECS-based analysis of active nematic channel flow, and review experimental and theoretical work on flow reversals in active and passive fluids. In Section III, we use the symmetries of the 2D AN equations to carry out an equivariant bifurcation analysis in minimal flow units, tracing primary and secondary local and global bifurcations, and constructing a symmetry-organized catalog of dynamically relevant ECSs. In Section IV, we use this catalog to describe transient flow reversals in the preturbulent regime, identifying a small number of ECS families and their heteroclinic connections as the main mechanisms by which trajectories switch between opposite nearly unidirectional states. In Section V, we extend this picture into the active turbulent regime in two MFUs, characterize the chaotic and quasiperiodic sets that arise at high activity, and demonstrate that persistent two-way reversals are organized by a reduced ECS skeleton whose members can be related back to the preturbulent branches via their symmetry signatures.
@@ II.1 Governing Equations
The domain is taken to be a two-dimensional channel of size $0,\tilde{L}]\times[0,h]$, with $x\in[0,\tilde{L}$] the streamwise (horizontal) direction and $y\in[0,h]$ the wall-normal (vertical) direction. We impose no-slip boundary conditions on the velocity field $\mathbf{u}$ and perpendicular (homeotropic) anchoring on the nematic director $\mathbf{n}$ at the confining walls. We employ the nondimensionalized equations obtained in [27]:
@@ II.3 Flow Reversals in Active Nematics and Other Active Fluids
The Ph.D. thesis [39] reported experimental results on AN systems in long, narrow channels whose open ends were connected to an external chaotic AN reservoir. Persistent, nearly unidirectional flows were observed, and the direction of these uniaxial flow states (UNI) was selected at random via spontaneous symmetry breaking of the zero-flow (ZF) state, as expected. More importantly, spontaneous and repeated reversals were also observed between the left- and right-flowing UNI states, and the reversal frequency depended on the channel dimensions. The author used the channel-averaged streamwise velocity $\langle u\rangle$ as an order parameter to quantify these reversals. In these experiments, a typical reversal proceeded via an intermediate state: the UNI state with a parabolic velocity profile ($\langle u\rangle\approx\pm u_{0}$) first collapsed into a vortex-dominated configuration with negligible net flow ($\langle u\rangle\approx 0$), and then evolved into a UNI state whose mean flow was opposite in direction to the original one (see Fig. 1).
Figure 1: Velocity field snapshots from active nematic experiments of Doré [39] of a reversal event in a channel of size $500\mu m\times 100\mu m$. The colormap shows the speed (µm/s). The flow evolves from a left-flowing UNI-like state ($\langle u\rangle<0$) to a right-flowing UNI-like state ($\langle u\rangle>0$) via an intermediate vortex-dominated configuration ($\langle u\rangle\approx 0$). Reproduced with permission of the author.
@@
In polar active fluids, recent experimental, computational and analytical work [40] has argued that flow reversals are a precursor to active turbulence. The authors found that a dense suspension of swimming Bacillus subtilis confined to a cylindrical wells exhibit a single vortex steady flow at small radii. This state, which is analogous to the UNI in channel flow, loses stability as the radius is increased beyond a threshold. In the resulting oscillatory state, the vortex periodically reverses its direction. Using simulations and weakly nonlinear analysis of a Toner–Tu–Swift–Hohenberg model, it was shown that the onset of reversals is controlled by a global infinite-period bifurcation mediated by a resonant nonlinear interaction of the three lowest modes of the flow.
@@ II.4 Possible Generic Dynamical Mechanisms Underlying Flow Reversals
Beyond active fluids, flow reversals have been extensively observed and studied in various settings [41, 42, 43, 44]. A prominent example is that of turbulent Rayleigh-Benard convection (RBC) [45], in which large-scale convection (LSC) reversals between the clockwise and anticlockwise vortical states are robustly observed [46]. In addition to numerical simulations, such reversals have been studied using both determininistic and stochastic reduced-order dynamical models [47, 42].
@@ III.2 Symmetries in the active nematics equations
Figure 2: Snapshots of the velocity field for selected ECSs at representative activity numbers $\mathrm{R_{a}}=1.5$ and $\mathrm{R_{a}}=4.5$, showing their symmetry generators in the full-length channel, $L=\tilde{L}$. The colormap represents vorticity.
@@ III.4.1 Traveling waves emerge from UNI
Figure 5: Projection onto the reduced phase space ($\langle Q_{11}\rangle$, $\langle v^{2}\rangle$), showing the stable and unstable TWs before the SNIPER bifurcation, with a heteroclinic connection passing close to UNI. Here, $\langle.\rangle$ denotes channel average. Note the RPO that emerges after SNIPER has a shape similar to that of the vanished heteroclinic network.
@@ III.5.3 One-vortex standing waves via symmetry-preserving bifurcation
Figure 14: Two dancing-disclination-like POs (DPOs) that co-exist at $\mathrm{R_{a}}=1.5$, projected onto the reduced phase space ($\langle Q_{11}\rangle$$\langle v^{2}\rangle$). Both solutions lie in the $\langle u\rangle=0$ subspace (i.e., no net streamwise flow) due to the $\sigma_{x}\sigma_{y}$ symmetry.
@@ III.5.5 Secondary Bifurcations Associated with Higher Unstable Eigenvalues of LAN
So far, our analysis has been limited to bifurcations associated with the first unstable eigenspace of LAN. As the activity number is varied from $\mathrm{R}_{\mathrm{a}}\approx 0.85$ to $\mathrm{R}_{\mathrm{a}}\approx 4.5$, a total of eight (equal) pairs of real eigenvalues eventually cross the imaginary axis. The isotropy subgroups of these eight unstable 2D eigenspaces are listed in Table 4 in the Appendix. While the study of all bifurcations is outside the scope of this work, we briefly discuss the bifurcations from the second unstable eigenspace, since these solutions play an important role in the turbulent regime we analyze later in this paper.
@@
Figure 15: Visualization of POT2a at $\mathrm{R_{a}}=2.25$. (a) A snapshot of the nematic director and velocity fields, where its $\sigma_{x}\sigma_{y}\tau_{x}(L/2)$ symmetry is apparent. (b) Projection onto the reduced phase space: ($\langle Q_{11}\rangle$, $\langle v^{2}\rangle$). This PO lies in the $\langle u\rangle=0$ subspace (i.e., no net streamwise flow) due to the $\sigma_{x}\sigma_{y}$ symmetry.
@@ IV.2 Transient Flow Reversals at Higher Activity and the Path Toward Turbulence
The chaotic sets $\mathcal{X}$ and $\mathcal{Y}$ exhibit reversals between regions where the flow velocity has a mean value $\langle u\rangle$ that persistently oscillates about positive and negative values, respectively. In the following section, we analyze these attractors to elucidate the mechanisms responsible for flow reversals in the MFU at $\mathrm{R}_{\mathrm{a}}=4.5$.
@@ V Flow Reversal Mechanisms in Turbulence: Two-Way Persistent Reversals
We find that an arbitrary perturbation (i.e., with no symmetry) always ends up in $\mathcal{Y}$, indicating that in MFU A, $\mathcal{Y}$ is the only almost-global attractor while the other two sets $\mathcal{X}$ and $\mathcal{Z}$ are saddles. Motivated by the apparent two-fold translation (i.e., $\tau_{x}(L_{A}/2)$) symmetry of $\mathcal{X}$ itself, we verify that in fact, all initial conditions with $\tau_{x}(L/2)$ symmetry end up in $\mathcal{X}$ in MFU A. Hence, we also analyze dynamics in the smaller MFU B, where $\mathcal{Y}$ is eliminated due to lack of two-fold translation symmetry, and we find that the halved-version of $\mathcal{X}$ is the only almost-global attractor that attracts arbitrary initial conditions almost surely. The quasiperiodic $\mathcal{Z}$, while still present in MFU B, again only attracts a measure zero set of trajectories that have the additional translation symmetry. Table 1 summarizes this discussion and answers Q1 posed above. Now we turn to answering Q2 and Q3.
@@ V.2.1 Dominant ECSs Within $\mathcal{X}$
(b) Reduced phase space projection showing the segment turbulent trajectory $X(t)$ in the interval $29000\tau\leq t\leq 30000\tau$ (black) and the six main ECSs (and their symmetry copies). The color scheme for the ECSs is the same as in (a). The orange circle indicates the beginning of the segment ($t=29000\tau$) and the magenta square indicates the end of the segment ($t=30000\tau$).
(c) Reduced phase space portrait showing the six main ECSs (and their symmetry copies) shadowed by the turbulent trajectory, along with their symmetry-related copies under $\sigma_{y}$.
@@
Across the 2,501 points, 40 of the 75 ECSs (along with their symmetry copies) appear as the closest state at least once. To isolate the most dynamically significant structures, we focus on the ten most frequently encountered ECSs. Together, these ten ECSs account for approximately 62% of all occurrences (see Table 6 in Appendix). As anticipated, each of these ten ECSs are invariant under $\tau_{x}(L_{A}/2)$ symmetry.
Figure 22: Symmetry-reduced distance plot showing the closest encounters between the chaotic trajectory $X(t)$ and the ten most relevant ECSs and their symmetry copies. The ECS list on the left follows the same color scheme as in Fig. 21(a). The plot shows a clear pattern of repeated transitions (i.e., reversals) between the Group 1 ‘left flowing’ (blue) and ’right flowing’ (red) ECSs, via the Group 2 intermediate (green) ECSs.
These ten ECSs serve as a reduced representation that captures the essential skeleton of the attractor $\mathcal{X}$’s dynamics. Six of these (‘Group 2’, shown in green) are natural candidates for mediating transitions between the red and blue regions, i.e., for enabling reversals. Figure 22 shows the distance plot between the trajectory and these ECSs, highlighting the paths of reversals. The remaining four ECSs (‘Group 1’, shown in red/blue), along with those from the green category, are closely approximated by the trajectory. For example, in the interval $29000\tau,29200\tau]$, during the first reversal of the series (see also Fig. 21a), the distance plot shows that the trajectory first approaches the blue ECS group, with the mean velocity oscillating around $\langle u\rangle\approx-2$. It then nears some ECSs of the green group, where $\langle u\rangle\approx 0$, before finally transitioning to a state where the mean velocity oscillates around $\langle u\rangle\approx 2$ and closely follows the red ECS group.
@@ V.2.2 Shadowing in $\mathcal{X}$
Observable reconstruction: Additional indirect evidence for shadowing comes from reconstruction of observables along a turbulent trajectory using ECS data. Motivated by periodic orbit theory [59], several recent works [74, 75, 76] have proposed approximating inertial turbulent statistics using a collection of invariant solutions, such as unstable periodic and relative periodic orbits. Adopting this perspective, we approximate a physical observable $\mathcal{O}$ in the active nematic system along a turbulent trajectory $X(t)$ by a weighted superposition of ECS contributions,
$\mathcal{O}\big(X(t)\big)\approx\sum_{i}w_{i}(t)\,\mathcal{O}\big(X_{E_{i}}(s_{i}(t))\big),$
where $X_{E_{i}}(s)$ denotes the state along the $i$-th ECS, $s_{i}(t)$ is the phase that minimizes the distance to $X(t)$ at time $t$, and $w_{i}(t)$ represents the instantaneous weight assigned to that ECS. As a minimal realization, we define the weights using a nearest-ECS criterion: at each time $t$, the ECS closest (in phase space) to the turbulent state is assigned unit weight, while all others are set to zero. We use three spatially averaged observables: the kinetic energy, enstrophy, and Frank elastic energy. These quantities have been previously used to characterize scaling behavior and statistical properties of active turbulent flows [21, 77]. We define them as follows:
$E_{k}$ $=\frac{1}{2A}\int_{A}\left(U^{2}+V^{2}\right)\,dx\,dy,$ (15)
$\mathcal{E}$ $=\frac{1}{2A}\int_{A}\left(\partial_{y}U-\partial_{x}V\right)^{2}\,dx\,dy,$ (16)
$E_{\text{Frank}}$ $=\frac{1}{2A}\int_{A}\left(|\nabla Q_{AA}|^{2}+|\nabla Q_{AB}|^{2}\right)\,dx\,dy,$ (17)
where $A$ is the channel area.
Our approximation captures the trends of all three observables remarkably well, as illustrated by the $8\tau$ moving average time series in Fig. 25. The kinetic energy is reproduced with particularly high fidelity: the correlation between the turbulent signal and its ECS-based approximation is $0.991$, and their time averages are $4.946$ and $4.746$, respectively, corresponding to a relative error of about $4.1\%$.
Although enstrophy and Frank elastic energy are harder to reconstruct because they depend on spatial derivatives of the velocity field and nematic director field, the time series in Fig. 25 nevertheless reproduces their overall trends reasonably well. Local agreement improves markedly during intervals of stronger shadowing; for instance, around $t=29100\tau$ and $t=29800\tau$. In general, however, the ECS-based approximation tends to underestimate both enstrophy and Frank energy.
Quantitatively, enstrophy exhibits a correlation of $0.848$ between the turbulent trajectory and its ECS approximation, with temporal averages $\langle\mathcal{E}\rangle=0.436$ and $0.400$, corresponding to a relative error of $8.3\%$. The Frank energy shows a correlation of $0.656$, with temporal averages $\langle E_{\mathrm{Frank}}\rangle=0.02896$ and $0.0278$, yielding a relative error of approximately $4.0\%$. These results indicate that even observables sensitive to finer spatial structures are well captured, providing further evidence that ECSs form the backbone of active turbulence in the minimal flow unit.
Figure 25: Moving averages (with window $8\tau$) of each observable, with solid curves showing the turbulent time series and dashed curves the corresponding ECS-based reconstructions.
@@ V.2.3 Reversal Mechanisms and Heteroclinic connections
First, the Group 2 ECSs play a role analogous to that of the ‘intermediary’ DPOs in the preturbulent regime. These ECSs are likely originated from solutions invariant under symmetries conjugate to $\sigma_{x}\sigma_{y}$ and $\tau_{x}(L_{A}/2)$ in the preturbulent regime. These ECSs in the turbulent regime have slightly broken $\sigma_{x}\sigma_{y}$ symmetry, but retain the two-fold $\tau_{x}(L_{A}/2)$ translational symmetry. Supporting evidence includes their similarity to POT2a in the reduced phase space. Second, the Group 1 ECSs play a role analogous to that of HRPOs, and form the structural backbone for generating left- and right-flowing, nearly uniaxial states. Their phase space structures suggests that these ECSs are most likely the result of symmetry-breaking bifurcations of the HRPOs. The Group 1 ECS retain the homoclinic-like structure in phase space, have broken the shift-reflect symmetries of their preturbulent counterparts, but still possess the two-fold $\tau_{x}(L_{A}/2)$ translation symmetry.
To move beyond qualitative resemblance, we quantify residual symmetry in the turbulent ECSs by measuring, along each ECS orbit $X_{E}(t)$, distances to the appropriate symmetry-related images $\tau_{x}(\ell)\sigma_{x}X_{E}(t)$ and $\tau_{x}(-\ell)\sigma_{x}\sigma_{y}\tau_{x}(\ell)X_{E}(t)$, for $\ell\in[0,L]$. Since both group 1 and 2 ECSs have two-fold translation symmetry, we work in the half domain MFU B with $L=L_{B}$. For each snapshot we minimize over $l$, and use the minimizing shift and the corresponding distance as a remanent symmetry signature, see Fig. 28. We find that Group 1 ECSs come markedly closer to their $l=L/2$ shift-reflect images, consistent with their descent from the unstable UNI eigenspace. Group 2 ECSs instead remain closer to images under the conjugates of $\sigma_{x}\sigma_{y}$, supporting their interpretation as descendants of LAN-related branches.
@@
Figure 28: Remanent symmetry analysis based on turbulent ECS orbit data in MFU B. Top: Minimum, mean, and median distances between snapshots along each ECS orbit and their symmetry-related images under the two symmetry types. The results indicate that the dominant approximate (remanent) symmetry is shift–reflect ($\sigma_{x}\tau_{x}(L_{B}/2)$) for Group 1, whereas for Group 2 it is a symmetry conjugate to the reflection $\sigma_{x}\sigma_{y}$. Bottom: Distributions of the optimal shifts $\ell$ for Group 1 ECSs, showing strong peaks near $L_{B}/2\approx 6.667$.
The heteroclinic connections between these groups are the likely mechanisms through which turbulence exhibits flow reversals. Indeed, we observe that heteroclinic connections may exist between pairs of ECSs from different regions that are shadowed sequentially. For example, the red ECS $\sigma_{x}\sigma_{y}$T2/1 is shadowed during the interval $20067.20\tau,\,29073.6\tau]$, and immediately afterward, the green ECS $\sigma_{y}$T2/30 is shadowed during $29072.80\tau,\,29073.6\tau]$. This suggests a heteroclinic connection between these two ECSs. However, within the green region, the turbulent trajectory sequentially (and sometimes simultaneously) shadows several ECSs, including $\sigma_{y}$T2/30, $\sigma_{x}$T2/25, T2/13, $\sigma_{y}$T2/9, $\sigma_{x}$T2/30, $\sigma_{x}\sigma_{y}$T2/4, and $\sigma_{x}\sigma_{y}$T2/13. Consequently, identifying a heteroclinic connection to a target green ECS is challenging, as shadowing events at the red/green and blue/green boundaries involve multiple ECSs and are not sharply defined. Although we cannot conclusively verify a heteroclinic connection between $\sigma_{x}\sigma_{y}$T2/1 and $\sigma_{y}$T2/30, the turbulent trajectory shadows the unstable manifold of the former ECS and eventually approaches the latter. Throughout the subsequent exploration of the green region, the trajectory continues to follow the unstable manifold, as illustrated by the segment of $X(t)$ and a nearby trajectory in Fig. 29 within the interval $29064.0\tau,\,29094.4\tau]$.
@@
In this MFU B based calculation, we focus on the two most relevant Group 1 (blue/red) ECSs and the three most relevant Group 2 (green) ECSs (and their symmetry counterparts). In this smaller domain, the RPOs T2/1 and T2/15 each have two unstable directions, while the RPOs T2/9, T2/30, and T2/4 have three. We consider the dominant unstable direction to generate a sample of trajectories within each unstable manifold. Several of these trajectories pass nearby and shadow other ECSs, providing evidence of heteroclinic connections. Table 2 lists few of the pairs of ECSs connected by such trajectories, together with the corresponding minimum distances and shadowing time intervals. The Figs. 30a and 30b highlight two such examples.